OpenStudy (anonymous):

What is -4 1/2 x (-6 2/3)? Kind of confusing, I know.

OpenStudy (anonymous):

\[-4\frac{ 1 }{ 2 }\times 6\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

but 6 and 2/3 is in parentheses.